# Stateful Large Language Model Serving with Pensieve

### v1

@@ 4. System Design
In this section, we describe the system design of Pensieve. At a high level, we aim to save a conversation’s KV-tokens across multiple turns, if there is enough cache space to do so. To realize the potential performance benefits, we need to make cache swapping and missing token recomputation fast and efficient, by developing a few techniques to address the challenges in §3.
Figure 5. Overview of Pensieve
@@ 4.1. System Overview
Figure 5 shows the overall system architecture. Pensieve consists of a single scheduler and a worker engine to manage the GPU. The scheduler has two jobs: 1) it is responsible for batching requests for execution, and 2) it ensures that requests in the batch have sufficient GPU memory for execution. For 1), the scheduler performs fine-grained iteration-level batching (batchmaker, ; orca, ; vllm, ) so that a new request can join the batch with existing requests while the latter are in the process of performing autoregressive generation. For 2), the scheduler manages the allocation of slots in the KV cache and when to swap between the GPU and CPU KV cache.
The worker also has two jobs: 1) it invokes kernels to execute the LLM computation for a batch of requests. 2) it performs the actual data movements between the GPU and CPU KV cache based on a batch’s cache plan received from the scheduler.
@@ 4.2. A unified batch scheduler
Pensieve performs fine-grained iteration-level batching (batchmaker, ; orca, ; vllm, ). But our batching strategy differs from existing LLM serving systems. ORCA only batches non-attention operations and handles attention operations for each request individually (orca, ). By contrast, Pensieve batches both non-attention and attention operations, like vLLM (vllm, ). However, unlike vLLM which only forms a batch among requests in the same phase and thus processes each of the two types of batches in a separate model execution, Pensieve handles both the prefill phase and generation phase in a unified way. In other words, the Pensieve scheduler forms a batch of requests regardless of which phase they are in. In the same batch, some could be in their generation phase while others are in their prefill phase. Such unified batching is made possible by Pensieve’s multi-query attention kernel design (§4.4.1).
@@
For unified batch formation, the scheduler concatenates the tokens to be processed from all requests. For each new request that has just joined the batch, it has a sequence of tokens corresponding to its prompt. For each existing request in the batch, it has a single token corresponding to the last generation step’s output token. By combining the prefill phase together with the generation phase, we avoid running separate small kernels and can thus improve GPU utilization.
@@ 4.3. KV cache management
Traditionally, the KV cache in GPU only serves as computation workspace. In Pensieve, the GPU KV cache also serves as storage space to cache the KV-tokens of recently completed requests of active conversations. Pensieve adopts a two-tier hierarchical caching strategy and uses the much larger CPU memory as the second tier cache space. The scheduler determines the caching plan and instructs the worker to perform the actual data movements between GPU and CPU.
@@
The scheduler tracks the amount of free GPU KV cache slots left. Before handing off a batch of requests to the worker, the scheduler tries to ensure that any new request’s past KV-tokens will reside in the GPU and there is sufficient GPU memory for the execution. Specifically, if a request in the batch has any KV-tokens that have been swapped out to the CPU memory or dropped, the scheduler determines how much additional GPU KV cache slots are needed to swap in or re-calculate those absent KV-tokens. If there is sufficient space, the scheduler instructs the worker to perform the necessary allocation and to start swapping those KV-tokens from the CPU as part of the batch’s cache plan.
Pipelined KV cache recovery.
The scheduler does not wait for a request’s KV-tokens to be fully swapped in from the CPU before handing it off to the worker for execution. Rather, we follow the pipelined approach (pipeswitch, ) to overlap computation with data transfer. Specifically, we exploit the fact that an LLM model has many layers and each layer’s KV-token is only used in this layer’s self attention calculation. Instead of waiting for all layers to finish data transfer before starting the execution, we initiate the transfer layer by layer and start model computation at the same time. The worker uses GPU events to preserve data dependency: it only starts a layer’s self-attention kernel once that layer’s KV-tokens have been fully copied to the GPU. Pipelining transfer with computation allows us to hide the swap-in latency.
Ahead-of-the-time swapping and the eviction policy.
@@
Since Pensieve tries to preserve KV-tokens in the GPU for reuse across multi-turn conversations, the scheduler does not immediately release a request’s GPU cache slots as soon as it finishes, unlike existing systems (vllm, ). Instead of waiting until the GPU cache has run out, the scheduler asks the worker to copy (aka swap out) selected KV-tokens to the CPU if less than a threshold (e.g. 50%) of the GPU cache slots are available. The corresponding GPU memory is not immediately released until the scheduler later decides to allocate the same slots to another conversation.
The scheduler selects which conversation to swap out by balancing two factors: 1) the number of KV-tokens the conversation has in the GPU cache. The larger the size is, the more free space can be released. 2) the amount of time since the conversation has last sent a request, as also used in LRU. The scheduler weighs both factors by multiplying them to get a score for each conversation and preferentially swaps out those with higher scores. Our swapping is fine-grained; the scheduler only evicts half of a conversation’s KV-tokens at a time, specifically the trailing half. The scheduler iteratively decides which conversation to swap out based on their scores until the GPU cache is no longer under memory pressure. The scheduler uses the same eviction policy for the CPU cache. Tokens evicted from the CPU are simply dropped.
Handling dropped tokens.
If a request has some of its KV-tokens dropped due to CPU memory pressure, we resort to recomputation to handle such dropped tokens. Figure 6 illustrates the most general scenario for the whereabouts of a new request’s KV-tokens. As we always evict cached tokens from the trailing end of a conversation, the GPU cache generally holds the request’s early tokens, and the CPU cache holds the middle, while the rest may have be dropped. The scheduler swaps in those tokens cached in the CPU. For dropped tokens, the scheduler will fetch their corresponding raw text tokens from the conversation history saved in a persistent store (Figure 5). These retrieved raw tokens will be merged into (i.e. prepended to) the new request’s prompt and become part of the batch’s input tokens, as shown in Figure 6. Thus, the dropped tokens will be recomputed as part of the prefill phase for the new request. The policy of evicting/dropping from a context’s trailing end makes such merging possible.
Figure 6. Recompute KV-tokens dropped by the CPU cache as part of the new request’s prefill phase.
Suspending requests during generation
@@
In spite of ahead-of-time eviction, the scheduler might still encounter scenarios when the generation phase runs out of GPU cache since requests’ decoding lengths are not known a priori. In this situation, the scheduler suspends some requests’ execution by taking them out of the current batch, swaps out their corresponding KV-tokens to the CPU and puts them back in the waiting queue. It chooses which request to suspend according to the descending order of their arrival time. As suspension causes increased latency (due to waiting for the swap-out), we try to avoid it by conservatively reserving 10% of GPU cache slots for the execution of existing requests that are in the generation phase. In other words, the scheduler stops adding new request into the running batch unless there are more than 10% free GPU cache slots.
@@
Figure 7. Single-query attention vs multi-query attention, softmax and rescaling omiited. $d$ is dimension of the embedding vector. red shaded area in multi-query attention score represents causal masking. $K$ and $V$ are logically contiguous, but physically placed at random locations when paged memory management is used.
4.4. Multi-query attention for non-contiguous cache
@@
How to combine existing KV-tokens in GPU cache with those just swapped in from the CPU? The most naive solution is to allocate a contiguous memory region in the GPU to hold them both. However, doing so would incur expensive memory copying, since KV-tokens are large. A more promising solution is to allocate separate space only for those swapped-in KV-tokens and to design an attention kernel implementation that can handle non-contiguous memory in its KV cache.
@@
vLLM’s PagedAttention kernel can handle non-contiguous KV cache in the generation phase (vllm, ). However, for the prefill phase, it still uses existing kernels which require all KV-tokens to reside in contiguous memory. We cannot use PagedAttention because it assumes each request in a batch has exactly one input token, which is the case for generation but not prefill. Thus, we refer to PagedAttention as a single-query attention kernel because it computes the attention scores between a single input token’s query representation (Q) and the KV cache. We need to build a multi-query attention kernel which supports performing attention between the the query representations (Q) of multiple input tokens per request and the KV cache over non-contiguous memory.
@@
Our new kernel is similar to that of vLLM to the extent that both kernels need to support loading KV cache from non-contiguous GPU global memory into on-chip shared memory. Their main difference is illustrated in Figure 7. As vLLM’s kernel performs attention between a request’s single input token and the existing KV-tokens, its underlying computation can be described as two matrix-vector multiplication operations, as shown in Figure 7  [^1] The batched version of PagedAttention performs batched matrix-vector multiplications.. In contrast, our kernel handles multiple input tokens for each request, computing attention scores between all pairs of input tokens and the conversation’s existing KV-tokens. Therefore, its underlying computation can be described as two matrix-matrix multiplication operations, and the batched version of our kernel performs batched matrix-matrix multiplications. Like vLLM, we fuse the two multiplication operations according to (flash, ). Because of the additional dimension in the Q tensor, our kernel has more parallelization and tiling opportunities on the GPU. However, care must be taken to handle the new challenge that different request in the batch have different number of input tokens.
@@
When multiple input tokens of a request are handled together in a kernel, we need to apply causal masking so that a later input token does not “attend” to the earlier tokens. This requires setting a corresponding upper triangular part of the attention score matrix to 0, as shown by the red shaded area in Figure 7. Causal masking is not needed for the single-query attention kernel since it only processes one input token. We fuse the causal masking operation inside the multi-query attention kernel to avoid materializing the intermediate attention score matrix (flash, ).
@@
Matrix-matrix multiplication kernels are much more complex than those for matrix-vector multiplication because they typically use sophisticated algorithms to extract additional data reuse opportunities with GPU’s on-chip shared memory and to leverage GPU’s tensor core primitives. Therefore, instead of trying to extend vLLM’s PagedAttention kernel for multi-query attention, we base our implementation on an existing multi-query kernel from PyTorch and extend it to handle non-contiguous KV cache. Our kernel uses the high performance thread-block level matrix-matrix multiplication provided by NVIDIA Cutlass template library (cutlass, ). Unlike PagedAttention, our kernel does not group multiple KV-tokens into blocks. Instead, each KV-token can potentially reside in a different memory location. Our kernel accepts a list of slot indices to gather KV-tokens from scattered memory locations. Such flexibility helps reduce internal fragmentation compared to the paged design.
@@ 4.4.1. Unifying the prefill and generation phase
As discussed in §4.2, in Pensieve’s batch formation, new requests in their prefill phase can be grouped together with existing requests in their generation phase. This is enabled by our multi-query attention kernel, because single-query attention performed in the generation phase can be treated as a special case of the multi-query attention with query size equal to 1.
More concretely, Pensieve’s scheduler concatenates the input tokens to be processed from all requests, regardless of whether they correspond to prompts of new requests or to last step’s output token from existing requests. Some auxiliary data structure is maintained to keep track of each request’s corresponding region. During execution by the worker, these batched input tokens are fed through linear layers to generate each token’s QKV representations. Newly generated KV-tokens are stored into the allocated slots in the GPU’s KV cache. Then the worker applies our multi-query attention kernel to produce the output tokens for all requests.

### v2

@@ 4. System Design
At a high level, we aim to save a conversation’s KV-tokens across multiple turns in a multi-tier GPU-CPU cache. To realize the potential performance benefits, we need to make cache swapping and dropped token recomputation efficient, by developing techniques to address the challenges in §3.
@@ 4.1. System Overview
Figure 7. System architecture of Pensieve
Figure 7 shows the system architecture. Pensieve consists of a single scheduler and multiple workers, each of which manages one GPU. The scheduler has two jobs: 1) it is responsible for batching requests for execution, and 2) it ensures that requests in the batch have sufficient GPU memory for execution. For 1), the scheduler performs fine-grained iteration-level batching (Gao et al., 2018; Yu et al., 2022; Kwon et al., 2023) so that a new request can join the batch with existing requests while the latter is in the process of performing autoregressive generation. For 2), the scheduler manages the allocation of slots in the KV cache and determines when to swap between the GPU and CPU KV cache. Each worker also has two jobs: 1) it invokes GPU kernels to process a batch of requests. 2) it performs the actual data movements between the GPU and CPU KV cache based on the batch’s cache plan as determined by the scheduler.
@@ 4.2. A unified batch scheduler
Pensieve performs fine-grained iteration-level batching (Gao et al., 2018; Yu et al., 2022; Kwon et al., 2023). However, our batching strategy differs from existing LLM serving systems. While Orca only batches non-attention operations and handles attention operations individually for each request (Yu et al., 2022), Pensieve and vLLM (Kwon et al., 2023) batch requests for both non-attention and attention operations. However, unlike vLLM which only forms a batch among requests in the same phase and thus processes each of the two types of batches in separate kernel invocation, Pensieve handles both the prefill phase and generation phase in a unified way. In other words, the Pensieve scheduler forms a batch of requests regardless of which phase they are in. In the same batch, some could be in their generation phase while others are in their prefill phase. Such unified batching is made possible by Pensieve’s multi-token attention kernel design (§4.4.1).
@@
For unified batch formation, the scheduler concatenates the tokens to be processed from all requests. For each new request joining the batch, its tokens include those corresponding to the request’s new user prompt. For each existing request in the batch, its token includes the one generated by the last generation step. By combining the prefill phase together with the generation phase, we avoid running separate small kernels and can thus improve GPU utilization.
@@ 4.3. KV cache management
Traditionally, the KV cache in GPU only serves as a computation workspace. In Pensieve, the GPU KV cache also serves as storage space to cache the KV-tokens of recently completed requests of active conversations. Pensieve adopts a two-tier hierarchical caching strategy and uses the much larger CPU memory as the second-tier cache space. The scheduler determines the caching plan and instructs the worker to perform the actual data movements between GPU and CPU.
@@
The scheduler tracks the amount of free GPU KV cache slots left. Before handing off a batch of requests to the worker, the scheduler tries to ensure that any new request’s past KV-tokens will reside in the GPU and there is sufficient GPU memory for the execution. Specifically, if a request in the batch has any KV-tokens that have been swapped out to the CPU memory or dropped, the scheduler determines how many additional GPU KV cache slots are needed to swap in or re-calculate those absent KV-tokens. If there is sufficient space, the scheduler instructs the worker to perform the necessary allocation and to start swapping those KV-tokens from the CPU as part of the batch’s cache plan.
4.3.1. Eviction Policy
Pensieve performs fine-grained token-level cache eviction and dropping. We design the eviction policy to express two kinds of preferences: 1) it preferentially evicts from older conversations, aka those that have been inactive for a longer period of time. This is based on the same LRU assumption that the least recently active conversation will not see activity for a longer time. 2) it preferentially evicts tokens from the leading end of a conversation’s history context. This is based on the observation, previously shown in Figure 4, that leading tokens of a conversation are cheaper to recompute than trailing ones. Below, we describe how our policy evicts according to these preferences.
Eviction granularity.
In order to reduce the overhead caused by frequent eviction decision-making and moving small amounts of memory over the PCIe bus, we group KV-tokens into chunks and make eviction decisions at the granularity of chunks. The chunk size is configurable. In our experiments, we find that setting the chunk size to 32 tokens works well.
The retention value of a chunk.
In order to combine both the LRU preference and the evicting from the front preference, we calculate a score for each chunk to capture its retention value. Specifically, the retention value of a chunk is $V=\frac{Cost(s,l)}{T}$, where $Cost(s,l)$ represents the cost of recomputing a chunk of size $s$ with a context of size $l$, and the denominator $T$ is the amount of time since the conversation was last active. Pensieve evicts chunks according to the ascending order of their retention values so that chunks with lower recomputation cost and/or those belonging to conversations with longer inactive periods are preferentially evicted.
Estimating the recomputation cost.
In order to calculate the retention value of a chunk, we need to estimate its recomputation cost. In particular, we view the cost to recompute the embedding of a chunk as the sum of recomputing the LLM model’s attention operation and the rest of the non-attention operation: $Cost(s,l)=Cost_{attention}(s,l)+Cost_{other}(s)$, where $s$ is the chunk size and $l$ is the size of the context to which the chunk “attends” for the attention operation. The cost of non-attention computation cost ($Cost_{other}(s)$) consists of linear layers, layer normalization, non-linear activation, etc., and therefore is independent of the context size. On the contrary, attention requires accessing and performing computation with all $l$ context tokens. Since the eviction decisions are made for fixed-size chunks of 32 tokens, we can simplify the cost function to become $Cost(l)=Cost_{attention}(l)+c$ where $Cost_{attention}(l)$ is the cost of performing attention operation for a chunk of 32 tokens with context length $l$, and $c$ is a constant capturing the cost of non-attention computation. We perform offline profiling to estimate $c$ as well as $Cost_{attention}(l)$ with varying context sizes. Since it’s not feasible to profile all possible context sizes, we profile context sizes that are powers of 2 and use the measured values to interpolate the cost for other context sizes.
4.3.2. Ahead-of-the-time swapping
@@
Since Pensieve tries to preserve KV-tokens in the GPU for reuse by a later request in the same conversation, the scheduler does not immediately release a request’s GPU cache slots as soon as it finishes, unlike existing systems (Kwon et al., 2023). Instead of waiting until the GPU cache has run out, the scheduler asks the worker to copy (aka swap out) selected KV-tokens to the CPU if less than a threshold (e.g. 25%) of the GPU cache slots are available. The corresponding GPU memory is reclaimed in a lazy manner and is not immediately released until the scheduler later decides to allocate the same slots to another conversation.
When the CPU cache runs out of space, the same eviction policy (§4.3.1) is used to decide which KV-tokens to drop. Performing ahead-of-the-time swapping allows the scheduler to overlap cache eviction with GPU computation, thus fully hiding the latency of swapping.
4.3.3. Pipelined KV cache recovery.
The scheduler does not wait for a request’s KV-tokens to be fully swapped in from the CPU before handing it off to the worker for execution. Rather, we follow the pipelined approach (Bai et al., 2020) to overlap computation with data transfer. Specifically, we exploit the fact that an LLM model has many layers and each layer’s KV-token is only used in this layer’s self-attention calculation. Instead of waiting for all layers to finish data transfer before starting the execution, we initiate the transfer layer by layer and start model computation at the same time. The worker uses GPU events to preserve data dependency: it only starts a layer’s self-attention kernel once that layer’s KV-tokens have been fully copied to the GPU. Pipelining transfer with computation allows us to hide the swap-in latency.
4.3.4. Handling dropped tokens.
Figure 8. How Pensieve handles dropped tokens. (a) concatenate tokens of the dropped context and new prompt as the input. (b) Calculate projections $Q$, $K$, $V$ (c) Update KV cache using newly calculated $K$, $V$; swap in KV-tokens previously evicted to the CPU. (d) Do attention for a batch of two sub-requests: one for the dropped context to attend to itself, and the other for the new prompt to attend to the entire context.
If a request has some of its KV-tokens dropped due to CPU memory pressure, we resort to recomputation to handle such dropped tokens. Figure 5 shows the whereabouts of a typical new request’s KV-tokens. As we always evict cached tokens from the leading end of a conversation, the GPU cache generally holds the request’s latest tokens and the CPU cache holds the middle, while the earlier ones may have been dropped.
The scheduler swaps in those tokens cached in the CPU. For dropped tokens, the scheduler will fetch their corresponding raw text tokens from the conversation history saved in a persistent store (Figure 7). These retrieved raw tokens will be merged into (i.e. prepended to) the new request’s prompt and become part of the batch’s input tokens, as shown in Figure 8 (step a). In the prefill phase, the embedding of dropped tokens and new prompt tokens are concatenated together as they are processed by successive model layers. At each Transformer layer, the Query, Key, and Value tensors are computed (step b). Key and Value are stored in KV cache, and Pensieve maintains the KV locations for the entire conversation context including previously cached tokens (step c), which can then be used to perform attention.
However, as discussed in §3.2, the challenge that comes with dropping leading tokens is that tokens in query tensor correspond to two disconnected ranges in the context, while all existing attention kernels assume that the Query tensor region is consecutive. To address this, we treat these two ranges as two sub-requests that happen to share portions of the underlying context. Each row in Figure 8 (step d) represents the Query tensor and its corresponding KV context locations of each sub-request of the original request. As our multi-token attention GPU kernel design (§4.4) can support Query tensors of variable lengths for different requests in the batch and also accept non-contiguous KV-token locations, Pensieve only needs to update auxiliary data structures and no memory copy is incurred when processing the sub-requests.
4.3.5. Suspending requests during generation
@@
Despite ahead-of-time eviction, the scheduler might still encounter scenarios when the generation phase runs out of GPU cache since a request’s decoding length is not known a priori. In this situation, the scheduler suspends some requests’ execution by taking them out of the current batch, swaps out their corresponding KV-tokens to the CPU, and puts them back in the waiting queue. It chooses which request to suspend according to the descending order of their arrival time. As suspension causes increased latency (due to waiting for the swap-out), we try to avoid it by conservatively reserving 10% of GPU cache slots for the execution of existing requests that are in the generation phase. In other words, the scheduler stops adding new requests into the running batch unless there are more than 10% free GPU cache slots.
@@
Figure 9. Single-token attention vs multi-token attention, softmax and rescaling omitted. $d$ is the dimension of the embedding vector. The red shaded area in the multi-token attention score represents causal masking. $K$ and $V$ are logically contiguous, but physically placed at random locations when paged memory management is used.
4.4. Multi-token attention for non-contiguous cache
@@
How to combine existing KV-tokens in the GPU cache with those just swapped in from the CPU? The most naive solution is to allocate a contiguous memory region in the GPU to hold them both. However, doing so would incur expensive memory copying, since KV-tokens are large. A more promising solution is to allocate separate space only for those swapped-in KV-tokens and to design an attention kernel implementation that can handle non-contiguous memory in its KV cache.
@@
vLLM’s PagedAttention kernel can handle non-contiguous KV cache in the generation phase (Kwon et al., 2023). However, for the prefill phase, it still uses existing kernels which require all KV-tokens to reside in contiguous memory. We cannot use PagedAttention because it assumes each request in a batch has exactly one input token, which is the case for generation but not prefill. Thus, we refer to PagedAttention as a single-token attention kernel because it computes the attention scores between a single input token’s query representation (Q) and the KV cache. We need to build a multi-token attention kernel that supports performing attention between the query representations (Q) of multiple input tokens per request and the KV cache over non-contiguous memory.
@@
Our new kernel is similar to that of vLLM to the extent that both kernels need to support loading KV cache from non-contiguous GPU global memory into on-chip shared memory. Their main difference is illustrated in Figure 9. As vLLM’s kernel performs attention between a request’s single input token and the existing KV-tokens, its underlying computation can be described as two matrix-vector multiplication operations, as shown in Figure 9  [^1] The batched version of PagedAttention performs batched matrix-vector multiplications.. In contrast, our kernel handles multiple input tokens for each request, computing attention scores between all pairs of input tokens and the conversation’s existing KV-tokens. Therefore, its underlying computation can be described as two matrix-matrix multiplication operations, and the batched version of our kernel performs batched matrix-matrix multiplications. Like vLLM, we fuse the two multiplication operations according to (Dao et al., 2022). Because of the additional dimension in the Q tensor, our kernel has more parallelization and tiling opportunities on the GPU. However, care must be taken to handle the new challenge that different requests in the batch have different numbers of input tokens.
@@
When multiple input tokens of a request are handled together in a kernel, we need to apply causal masking so that an earlier input token does not “attend” to later tokens. This requires setting a corresponding upper triangular part of the attention score matrix to 0, as shown by the red shaded area in Figure 9. Causal masking is not needed for the single-token attention kernel since it only processes one input token. We fuse the causal masking operation inside the multi-token attention kernel to avoid materializing the intermediate attention score matrix (Dao et al., 2022).
@@
Matrix-matrix multiplication kernels are much more complex than those for matrix-vector multiplication because they typically use sophisticated algorithms to extract additional data reuse opportunities with GPU’s on-chip shared memory and to leverage GPU’s tensor core primitives. Therefore, instead of trying to extend vLLM’s PagedAttention kernel for multi-token attention, we base our implementation on an existing multi-token attention kernel from PyTorch and extend it to handle non-contiguous KV cache. Our kernel uses the high-performance thread-block level matrix-matrix multiplication provided by NVIDIA Cutlass template library (NVIDIA, 2017b).
@@ 4.4.1. Unifying the prefill and generation phase
As discussed in §4.2, in Pensieve’s batch formation, new requests in their prefill phase can be grouped with existing requests in their generation phase. This is enabled by our multi-token attention kernel because single-token attention performed in the generation phase can be treated as a special case of multi-token attention with query size equal to 1.
More concretely, Pensieve’s scheduler concatenates the input tokens to be processed from all requests, regardless of whether they correspond to prompts of new requests or the last step’s output token from existing requests. Some auxiliary data structure is maintained to keep track of each request’s corresponding region. During execution by the worker, these batched input tokens are fed through linear layers to generate each token’s QKV representations. Newly generated KV-tokens are stored in the allocated slots in the GPU’s KV cache. Then the worker applies our multi-token attention kernel to produce the output tokens for all requests.